Calculate the rate the infilation rate of the economy if the consumer price index for 2019 is 100.


1
Expert's answer
2022-03-02T10:26:41-0500

The inflation rate is calculated as follows;


"=\frac{(Current \space CPI\space - \space Past \space CPI)}{Past \space CPI} {\times100}"


"=\frac{100-90}{100}\times100"


"=\frac{10}{100}\times100"


"=10 \%"